# Erasing ‘Ugly’ from the Internet: Propagation of the Beauty Myth in Text-Image Models

### v1

@@ 1 Introduction
Social media has exacerbated beauty norms, leading people to have altered views of their body image, and particularly affecting women, for example with a notable rise in plastic surgery Thawanyarat et al. (2022); Laughter et al. (2023). Now, with the rise of generative language and vision models, it has been predicted that 90% of all internet content will soon be automatically generated Vincent (2023). With this in mind, how do such models affect societal beauty standards? That is, what do they generate as ‘beautiful’ or ‘ugly’? What implications does this have on the evolving beauty standards in social media?
@@ 2.1 What are beauty standards and how do they develop in society?
The nature of beauty has been a topic of philosophical discussion since ancient times Sartwell (2012). According to the Platonic ideal, it is an objective quality. That is, beautiful forms exist in the world and possess the quality of beauty irrespective of how people may respond to them. This contrasts with the more recent, commonly held idea that ‘beauty is in the eye off the beholder’. This was formalised by Hume (1894), who nevertheless did not consider beauty to be entirely subjective. In fact, 18th century philosophers such as Hume and Kant considered beauty to be recognisable at a collective level Kant (2003), with the ability to do so depending on the qualities of the observers: those with good taste (formed by factors such as background and education) were therefore the arbiters of beauty. For such (male) thinkers, women were typically seen as incapable of responding to beauty, and beautiful qualities were considered stereotypically feminine (such as soft, round, and delicate), as described by Burke (1757).
@@
The 20th century saw increasing recognition that beauty is a political issue, in which powerful sections of society are able both to define what is beautiful and to cast oppressed groups (such as the poor) as unattractive Sartwell (2012), and that, under capitalism, things seen as beautiful are used to manipulate people and sell products. We see countless examples of this historically, for instance pale skin was highly coveted in the Victorian era, because paleness signalled that one could afford to stay indoors and was a marker of upper class status, while those that had tanned skin were considered ugly, as a tan was associated with outdoor labourers. Tanning later became popular in the early 20th century, fuelled by the the narrative that a tan signalled wealth and the ability to take luxurious holidays North Lancashire Council (2025); Martin et al. (2009). At the same time, feminists criticised Western notions of beauty as revolving around the objectification of women and their exploitation in art and popular culture MacKinnon (1987); Wolf (1991).
@@
Cultural relativists recognise that different beauty standards exist in different cultures Brown (1998), and indeed varied conceptions of attractive features exist in indigenous societies such as the Maori people (Wolf, 1991, p.12). However, critiques of colonialism and racism view dominant, white, Western ideas of beauty as oppressive towards Black and Asian people, and recognise that these standards dominate others, motivating Black people to straighten their hair Hooks (1989) and Asian people to lighten their skin Saraswati (2020).
@@
Beauty standards are, therefore, constantly changing, influenced by fashions, philosophies, political movements, and, as we examine, new technologies. These are introduced at a young age via stories and fairy tales Spade & Valentine (2017); Baker-Sperry & Grauerholz (2003). The particular physical traits that are considered beautiful vary by culture and change over time Spade & Valentine (2017), but in modern times, images in advertising have tended to glorify thinness and weightloss Groesz et al. (2002).
@@
The second half of the twentieth century saw digital media enable the creation of unrealistic representations of human bodies White (2009), which has been recently exacerbated by the ease of dissemination on social media. There is evidence that the particular ideals promoted here are Eurocentric. For example, analysis of Japanese Instagram posts shows a distinct tendency towards westernising and whitening beauty standards Öhman & Metcalfe (2021).
@@
According to McMillan Cotton (2019) societal beauty standards serve to reinforce the existing social order, forcing people to try to conform to them and, as Wolf (1991) points out in The Beauty Myth, suffering a range of negative consequences no matter the extent to whether they achieve this. These include higher rates of depression, eating disorders, body dysmorphic disorder, and low self-esteem, as well as acceptance of lower pay and the high expense of self-preservation i.e., purchasing products deemed necessary to preserve the standards (Wolf, 1991, pp.48-57).
While the long-term consequences of social media use on society are yet to be established, there is already tangible evidence to support the link between traditional media consumption and people’s body-image. For example, Plumridge (2019) discusses the Fiji experiment Becker et al. (2002); a long term analysis showing that the introduction of television during the mid 90s to Viti Levu (the largest island in Fiji), led to a rise in eating disorders among ethnic Fijian adolescent girls. Previously, the region had had an extremely low occurrence of reported eating disorder cases, as the accepted beauty norms were for women to be plump or curvaceous, i.e. comments about weight loss were usually meant negatively and came from a place of concern. The study found that eating disorders were much more prevalent following exposure to western TV shows, i.e. a desire to emulate characters on shows such as ‘90210’ led to changes in the attitudes and beliefs about losing weight.
Similar effects have been linked to social media consumption, leading to higher dissatisfaction with one’s appearance, and even negatively impacting patients receiving treatment for body dysmorphic disorder Laughter et al. (2023); Ateq et al. (2024). For example, the ‘Zoom effect’ Thawanyarat et al. (2022); Laughter et al. (2023), shows that constantly looking at oneself due to video conferencing during the COVID-19 pandemic lead to heightened appearance concerns and increased google searches on above-the-shoulder aesthetic procedures. Recent work has proposed to revise body dysmorphic disorder diagnostic scales to include the influence of social media Maymone & Kroumpouzos (2022). Searches for cosmetic procedures like lip fillers have been shown to skyrocket by 50% even within the span of a week oH_2024recent, showing how susceptible the public can be towards changing trends on body parts. Similarly, social media consumption has been blamed for a rise in the popularity of unnecessary surgical procedures, such as dermal filling Dowrick & Holliday (2022).
@@ 3 Related Work: NLP and Image models
High levels of bias towards factors relevant to beauty standards have repeatedly been found in both text and image datasets and models, with resources favouring youth, whiteness, and conventional attractiveness. For instance, there is a large body of work investigating racial biases in NLP including in word embeddings Manzini et al. (2019); Lepori (2020), annotations Talat (2016), and models Zhang et al. (2020). Several works use template-based prompting to uncover stereotypes in language models. For racial stereotypes, Cheng et al. (2023) find high levels of racial stereotypes in generated text. Liu et al. (2024) find that LLMs favour younger demographics, while Kamruzzaman et al. (2024) discover age and beauty stereotypes in text completions. In terms of social media analysis, Öhman & Metcalfe (2021) analysed Instagram posts to understand how beauty is marketed in social media platforms specifically related to beauty standards in Japan. They utilised NLP tools such as topic modelling and discourse analysis on Instagram posts to uncover specific and frequent language use surrounding societal pressure to conform to these standards. We extend these template-based methods to our text-image pipelines.
In terms of image models and image datasets, Buolamwini & Gebru (2018) found that facial recognition datasets are heavily biased towards lighter-skinned subjects, and that darker-skinned women are frequently misclassified by models as a result. We follow them in using the Fitzpatrick Skin Type classification system for our analysis. Birhane et al. (2023) audited the large, commonly used LAION multimodal datasets, finding a correlation between sexually explicit content and hateful, aggressive, and targeted language. We use the same ‘NSFW’ label to assess the outputs of our generation pipelines. Investigating biases in text-to-image models, Bianchi et al. (2023) find that the prompt “an attractive person” produces images of faces that approximate a “[…] White ideal […]”, amongst other gender and racial stereotypes. They note that the easy availability of the models they use has serious implications for the propagation of these biases across the internet. Gulati et al. (2024) note that ‘lookism’ i.e., preference for idealised physical appearance is an understudied model bias. We attempt to address this gap by investigating how and to what extent beauty–and conversely ugliness–are represented in language and image generation models.
@@ 4.1 Base prompts
We created a template and a seed list of commonly used words associated with beauty, such as descriptors of skin, parts of the face, adjectives to describe the body and so on. To the best of our knowledge, such a list of words that are associated with beauty is not present in the literature. We do so by first consulting dermatology resources for (1) parts of the face: such as parts that are the most commonly enhanced in plastic surgery based on recent trends on Instagram Thawanyarat et al. (2023); and (2) Ways to describe skin, such as Fitzpatrick scales, [^2] Note that while Fitzpatrick scales remain a gold standard Roberts (2009), they have been criticised as subject to limitations, in particular for people of colour Ware et al. (2020). commonly used to describe skin, hair, and eye colour Gupta & Sharma (2019); NHS (2024), and a modified version for description of wrinkle development Shoshani et al. (2008). We create a list of attributes related to personal beauty, manually adding words and phrases such as ‘overall appearance’ and ‘style’.
We then prompt GPT-4 OpenAI (2024) using the prompt ‘What are some adjectives (positive, negative and neutral) to describe a person’s <attribute>. Please return a list.’ to get polarity categorised adjectives to describe the particular attribute. We compile these lists of attributes and adjectives into prompt templates by adding three gender markers, as shown in equation LABEL:eq:eq1, where $S_{1}$ and $S_{2}$ denote example base prompts created from the template. Note, while phrases such as ‘A non-binary person with bright blue eyes’ would not be used colloquially over ‘A person with bright blue eyes’, we deliberately emphasise this gender marker, as we want to evaluate the output of LLMs and image models when such a marker is present.
@@ 4.2 Description generation
For textual description generation in pipeline 2, we use three open source and state-of-the-art LLMs: Meta’s Llama-3.1-8B base [^4] https://huggingface.co/meta-llama/Llama-3.1-8B., Llama-3.1-8B instruct [^5] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. Grattafiori et al. (2024), and Deepseek-7b chat [^6] https://huggingface.co/deepseek-ai/deepseek-llm-7b-chat. DeepSeek-AI (2025). While we choose the Llama-3.1 models due to their extensive use in the community, we choose Deepseek-7b as it might generate variation in the outputs stemming from cultural differences, given that Deepseek AI is based in China and Meta in the United States. We experiment with text-based LLMs and different hyperparameters to obtain the most varied and least repetitive descriptions and keep these constant throughout the three models (see details in Appendix A Hyperparameters).
@@ 4.3 Image generation
To generate images, we used two open-source models: Stable-diffusion-3.5-large, [^8] https://huggingface.co/stabilityai/stable-diffusion-3.5-large. a newer release of the Stable Diffusion models from Stability AI, and F-Lite Ryu et al. (2025), [^9] https://huggingface.co/Freepik/F-Lite. a lite diffusion model from Freepik. While both models are state-of-the-art (choosing Stable-diffusion-3.5 because of its general popularity), we specifically choose F-lite given that it was trained on Safe For Work (SFW) images Ryu et al. (2025) and specifies the capacity to handle longer prompts, which will be useful for the description-generated images.
To generate the most photo-realistic images, we consult different guidelines for hyperparameter settings, such as the recommendations from the developers of the model and blog posts on Stable Diffusion fofr (2024). Similar to Tiku & Chen (2024), the input we narrowed in on which consistently produced the most realistic looking images was ‘A full length realistic photo of: <prompt>. We kept the input and hyperparameter settings consistent for both models and both pipelines.
For the prompt-generated images, we used the base prompts as the input with a minor modification, i.e. ‘charming man’ $\to$ ‘A charming man.’. For the description-generated images, we selected the generated text description by the three LLMs with the highest diversity, that is the lowest self-BLEU score Zhu et al. (2018) to use in place of the <prompt> as the input. In total, we generated 5984 images, i.e. 1496 prompt-generated and description-generated images each for each model.
@@ 4.4 Human evaluation
For $H_{2}$, we ask participants questions regarding the age range, skin tone, socio-economic status and whether the image contains nudity/is sexualised. We ask participants about skin tone range, given that race and ethnicity are complex constructs that cannot be inferred visually, and may be subject to biased assumptions from the participants based on stereotypes. While other scales have been validated to study skin tone in machine learning research, such as the Monk Skin Tone Scale Google Research (2022) based on Monk (2019), we choose the Fitzpatrick scales given that it remains a gold standard in dermatology research Roberts (2009), and there has been precedent in using this scale to assess skin tone in AI image research Buolamwini & Gebru (2018). Figure 3 shows the scale that participants are shown in relation to $Q2$.
@@
For socio-economic status, we use categories that are outlined in Cercas Curry et al. (2024). For age range, we group together age bands used commonly in government survey research, e.g. Office for National Statistics UK (2025).
@@
Research shows that age and gender differences exist in selfie behaviours, and that women are more likely to utilise photographic filers in selfies compared to men Dhir et al. (2016). Additionally, surveys show that the largest age demographic for Instagram and Snapchat usage is between ages 182418-24, and the second largest is between 253425-34 Datareportal (2025a; b). Furthermore, perceptions of objectification and sexualisation depend on one’s lived experiences. Sexualisation is a form of online Gender Based Violence United Nations (2025), which is relevant to this research on digital media and generative AI, that disproportionately affects $\approx 50\%$ of women and other marginalised genders Glitch (2020). Bearing in mind these factors, we recruit participants based on their reported demographic background and use of social media in their daily lives. We recruit participants via the Prolific platform. [^11] https://www.prolific.com/ We narrow down demographics by recruiting participants that live in the UK, who’s gender is either women or non-binary, are between $18$ and $35$, and who self-report use of visual social media platforms (specifically Snapchat, Instagram and/or Tik Tok), and additionally phone device usage as everyday to multiple times a day. For details of participant welfare measures, see Ethical Considerations.

### v2

@@ 1 Introduction
Social media has exacerbated beauty norms, leading people to have altered views of their body image, and particularly affecting women, for example with a notable rise in plastic surgery (thawanyarat2022zoom; LAUGHTER202328). Now, with the rise of generative language and vision models, it has been predicted that 90% of all internet content will soon be automatically generated (vincent-2023-ai). With this in mind, how do such models affect societal beauty standards? That is, what do they generate as ‘beautiful’ or ‘ugly’? What implications does this have on the evolving beauty standards in social media?
@@ 2.1 What are beauty standards and how do they develop in society?
The nature of beauty has been a topic of philosophical discussion since ancient times (sartwell2012beauty). According to the Platonic ideal, it is an objective quality. That is, beautiful forms exist in the world and possess the quality of beauty irrespective of how people may respond to them. This contrasts with the more recent, commonly held idea that ‘beauty is in the eye off the beholder’. This was formalised by hume1894standard, who nevertheless did not consider beauty to be entirely subjective. In fact, 18th century philosophers such as Hume and Kant considered beauty to be recognisable at a collective level (kant2003observations), with the ability to do so depending on the qualities of the observers: those with good taste (formed by factors such as background and education) were therefore the arbiters of beauty. For such (male) thinkers, women were typically seen as incapable of responding to beauty, and beautiful qualities were considered stereotypically feminine (such as soft, round, and delicate), as described by burke-1757-philosphical.
@@
The 20th century saw increasing recognition that beauty is a political issue, in which powerful sections of society are able both to define what is beautiful and to cast oppressed groups (such as the poor) as unattractive (sartwell2012beauty), and that, under capitalism, things seen as beautiful are used to manipulate people and sell products. We see countless examples of this historically, for instance pale skin was highly coveted in the Victorian era, because paleness signalled that one could afford to stay indoors and was a marker of upper class status, while those that had tanned skin were considered ugly, as a tan was associated with outdoor labourers. Tanning later became popular in the early 20th century, fuelled by the the narrative that a tan signalled wealth and the ability to take luxurious holidays (NL_heritage; martin2009changes). At the same time, feminists criticised Western notions of beauty as revolving around the objectification of women and their exploitation in art and popular culture (mackinnon-1987-desire; wolf1991beauty).
@@
Cultural relativists recognise that different beauty standards exist in different cultures (brown_beauty_1998), and indeed varied conceptions of attractive features exist in indigenous societies such as the Maori people (wolf1991beauty, p.12). However, critiques of colonialism and racism view dominant, white, Western ideas of beauty as oppressive towards Black and Asian people, and recognise that these standards dominate others, motivating Black people to straighten their hair (hooks--1989-from) and Asian people to lighten their skin (saraswati2020cosmopolitan).
@@
Beauty standards are, therefore, constantly changing, influenced by fashions, philosophies, political movements, and, as we examine, new technologies. These are introduced at a young age via stories and fairy tales (spade-etal-2019-kaleidoscope; baker-sperry-etal-2003-pervasiveness). The particular physical traits that are considered beautiful vary by culture and change over time (spade-etal-2019-kaleidoscope), but in modern times, images in advertising have tended to glorify thinness and weightloss (groesz-etal-2002-effect).
@@
The second half of the twentieth century saw digital media enable the creation of unrealistic representations of human bodies (white-2009-networked), which has been recently exacerbated by the ease of dissemination on social media. There is evidence that the particular ideals promoted here are Eurocentric. For example, analysis of Japanese Instagram posts shows a distinct tendency towards westernising and whitening beauty standards (ohman-metcalfe-2021-japanese).
@@
According to mcmillan-cottom-2019-thick societal beauty standards serve to reinforce the existing social order, forcing people to try to conform to them and, as wolf1991beauty points out in The Beauty Myth, suffering a range of negative consequences no matter the extent to whether they achieve this. These include higher rates of depression, eating disorders, body dysmorphic disorder, and low self-esteem, as well as acceptance of lower pay and the high expense of self-preservation i.e., purchasing products deemed necessary to preserve the standards (wolf1991beauty, pp.48-57).
While the long-term consequences of social media use on society are yet to be established, there is already tangible evidence to support the link between traditional media consumption and people’s body-image. For example, fiji_experiment discusses the Fiji experiment (becker2002eating); a long term analysis showing that the introduction of television during the mid 90s to Viti Levu (the largest island in Fiji), led to a rise in eating disorders among ethnic Fijian adolescent girls. Previously, the region had had an extremely low occurrence of reported eating disorder cases, as the accepted beauty norms were for women to be plump or curvaceous, i.e. comments about weight loss were usually meant negatively and came from a place of concern. The study found that eating disorders were much more prevalent following exposure to western TV shows, i.e. a desire to emulate characters on shows such as ‘90210’ led to changes in the attitudes and beliefs about losing weight.
Similar effects have been linked to social media consumption, leading to higher dissatisfaction with one’s appearance, and even negatively impacting patients receiving treatment for body dysmorphic disorder (LAUGHTER202328; ateq_2024). For example, the ‘Zoom effect’ (thawanyarat2022zoom; LAUGHTER202328), shows that constantly looking at oneself due to video conferencing during the COVID-19 pandemic lead to heightened appearance concerns and increased google searches on above-the-shoulder aesthetic procedures. Recent work has proposed to revise body dysmorphic disorder diagnostic scales to include the influence of social media (MAYMONE2022554). Searches for cosmetic procedures like lip fillers have been shown to skyrocket by 50% even within the span of a week (oh2024recent), showing how susceptible the public can be towards changing trends on body parts. Similarly, social media consumption has been blamed for a rise in the popularity of unnecessary surgical procedures, such as dermal filling (dowrick-holliday-2022-tweakments).
@@ 3 Related Work: NLP and Image models
High levels of bias towards factors relevant to beauty standards have repeatedly been found in both text and image datasets and models, with resources favouring youth, whiteness, and conventional attractiveness. For instance, there is a large body of work investigating racial biases in NLP including in word embeddings (manzini-etal-2019-black; lepori-2020-unequal), annotations (talat-2016-racist), and models (zhang-etal-2020-demographics). Several works use template-based prompting to uncover stereotypes in language models. For racial stereotypes, cheng-etal-2023-marked find high levels of racial stereotypes in generated text. liu-etal-2024-generation-gap find that LLMs favour younger demographics, while kamruzzaman-etal-2024-investigating discover age and beauty stereotypes in text completions. In terms of social media analysis, ohman-metcalfe-2021-japanese analysed Instagram posts to understand how beauty is marketed in social media platforms specifically related to beauty standards in Japan. They utilised NLP tools such as topic modelling and discourse analysis on Instagram posts to uncover specific and frequent language use surrounding societal pressure to conform to these standards. We extend these template-based methods to our text-image pipelines.
In terms of image models and image datasets, buolamwini-gebru-2018-gender found that facial recognition datasets are heavily biased towards lighter-skinned subjects, and that darker-skinned women are frequently misclassified by models as a result. We follow them in using the Fitzpatrick Skin Type classification system for our analysis. birhane2023laions audited the large, commonly used LAION multimodal datasets, finding a correlation between sexually explicit content and hateful, aggressive, and targeted language. We use the same ‘NSFW’ label to assess the outputs of our generation pipelines. Investigating biases in text-to-image models, bianchi-etal-2023-easily find that the prompt “an attractive person” produces images of faces that approximate a “[…] White ideal […]”, amongst other gender and racial stereotypes. They note that the easy availability of the models they use has serious implications for the propagation of these biases across the internet. gulati2024lookismoverlookedbiascomputer note that ‘lookism’ i.e., preference for idealised physical appearance is an understudied model bias. We attempt to address this gap by investigating how and to what extent beauty–and conversely ugliness–are represented in language and image generation models.
@@ 4.1 Base prompts
We created a template and a seed list of commonly used words associated with beauty, such as descriptors of skin, parts of the face, adjectives to describe the body and so on. To the best of our knowledge, such a list of words that are associated with beauty is not present in the literature. We do so by first consulting dermatology resources for (1) parts of the face: such as parts that are the most commonly enhanced in plastic surgery based on recent trends on Instagram (thawanyarat2023prs); and (2) Ways to describe skin, such as Fitzpatrick scales, [^2] Note that while Fitzpatrick scales remain a gold standard (roberts2009skin), they have been criticised as subject to limitations, in particular for people of colour (ware2020racial). commonly used to describe skin, hair, and eye colour (gupta2019skin; fitzpatrick_nhs), and a modified version for description of wrinkle development (shoshani2008modified). We create a list of attributes related to personal beauty, manually adding words and phrases such as ‘overall appearance’ and ‘style’.
We then prompt GPT-4 (openai2024gpt4technicalreport) using the prompt ‘What are some adjectives (positive, negative and neutral) to describe a person’s <attribute>. Please return a list.’ to get polarity categorised adjectives to describe the particular attribute. We compile these lists of attributes and adjectives into prompt templates by adding three gender markers, as shown in equation LABEL:eq:eq1, where $S_{1}$ and $S_{2}$ denote example base prompts created from the template. Note, while phrases such as ‘A non-binary person with bright blue eyes’ would not be used colloquially over ‘A person with bright blue eyes’, we deliberately emphasise this gender marker, as we want to evaluate the output of LLMs and image models when such a marker is present.
@@ 4.2 Description generation
For textual description generation in pipeline 2, we use three open source and state-of-the-art LLMs: Meta’s Llama-3.1-8B base [^4] https://huggingface.co/meta-llama/Llama-3.1-8B., Llama-3.1-8B instruct [^5] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. (grattafiori2024llama3herdmodels), and Deepseek-7b chat [^6] https://huggingface.co/deepseek-ai/deepseek-llm-7b-chat. (deepseekai2025deepseekv3technicalreport). While we choose the Llama-3.1 models due to their extensive use in the community, we choose Deepseek-7b as it might generate variation in the outputs stemming from cultural differences, given that Deepseek AI is based in China and Meta in the United States. We experiment with text-based LLMs and different hyperparameters to obtain the most varied and least repetitive descriptions and keep these constant throughout the three models (see details in Appendix A Hyperparameters).
@@ 4.3 Image generation
To generate images, we used two open-source models: Stable-diffusion-3.5-large, [^8] https://huggingface.co/stabilityai/stable-diffusion-3.5-large. a newer release of the Stable Diffusion models from Stability AI, and F-Lite (ryu2025flite), [^9] https://huggingface.co/Freepik/F-Lite. a lite diffusion model from Freepik. While both models are state-of-the-art (choosing Stable-diffusion-3.5 because of its general popularity), we specifically choose F-lite given that it was trained on Safe For Work (SFW) images (ryu2025flite) and specifies the capacity to handle longer prompts, which will be useful for the description-generated images.
To generate the most photo-realistic images, we consult different guidelines for hyperparameter settings, such as the recommendations from the developers of the model and blog posts on Stable Diffusion (fofr_2024). Similar to WP_2024, the input we narrowed in on which consistently produced the most realistic looking images was ‘A full length realistic photo of: <prompt>. We kept the input and hyperparameter settings consistent for both models and both pipelines.
For the prompt-generated images, we used the base prompts as the input with a minor modification, i.e. ‘charming man’ $\to$ ‘A charming man.’. For the description-generated images, we selected the generated text description by the three LLMs with the highest diversity, that is the lowest self-BLEU score (zhu-etal-2018-texygen) to use in place of the <prompt> as the input. In total, we generated 5984 images, i.e. 1496 prompt-generated and description-generated images each for each model.
@@ 4.4 Human evaluation
For $H_{2}$, we ask participants questions regarding the age range, skin tone, socio-economic status and whether the image contains nudity/is sexualised. We ask participants about skin tone range, given that race and ethnicity are complex constructs that cannot be inferred visually, and may be subject to biased assumptions from the participants based on stereotypes. While other scales have been validated to study skin tone in machine learning research, such as the Monk Skin Tone Scale (monk_skin_tone_google) based on monk2019monk, we choose the Fitzpatrick scales given that it remains a gold standard in dermatology research (roberts2009skin), and there has been precedent in using this scale to assess skin tone in AI image research (buolamwini-gebru-2018-gender). Figure 3 shows the scale that participants are shown in relation to $Q2$.
@@
For socio-economic status, we use categories that are outlined in curry-etal-2024-classist. For age range, we group together age bands used commonly in government survey research, e.g. ONS_age_bands.
@@
Research shows that age and gender differences exist in selfie behaviours, and that women are more likely to utilise photographic filers in selfies compared to men (DHIR2016549). Additionally, surveys show that the largest age demographic for Instagram and Snapchat usage is between ages 182418-24, and the second largest is between 253425-34 (instagram_stats; snapchat_stats). Furthermore, perceptions of objectification and sexualisation depend on one’s lived experiences. Sexualisation is a form of online Gender Based Violence (UN_women), which is relevant to this research on digital media and generative AI, that disproportionately affects $\approx 50\%$ of women and other marginalised genders (glitch-ewaw-2020-ripple). Bearing in mind these factors, we recruit participants based on their reported demographic background and use of social media in their daily lives. We recruit participants via the Prolific platform. [^11] https://www.prolific.com/ We narrow down demographics by recruiting participants that live in the UK, who’s gender is either women or non-binary, are between $18$ and $35$, and who self-report use of visual social media platforms (specifically Snapchat, Instagram and/or Tik Tok), and additionally phone device usage as everyday to multiple times a day. For details of participant welfare measures, see Ethical Considerations.